# Certifying nonlocal properties of noisy quantum operations

### v1

@@ V Conclusions and open questions
We infer nonlocal properties of bipartite quantum channels based on the bipartite stochastic matrices produced by them under measurement protocols. In particular, our approach allows us to bound the sets of local operations assisted by shared randomness (LOSR), local operations assisted by shared entanglement (LOSE) and quantum non-signaling channels (qNS), from the conditional probabilities of local measurement outcomes. This framework is suitable to experimental verification of nonlocality of bipartite channels from local measurements, even in the presence of complete decoherence. In contrast, we show that both entanglement and nonlocality of bipartite quantum channels can be enhanced under nonlocal gradual decoherence.
@@
Figure 7: First glimpse at the structure of multidimensional convex sets of quantum and classical: states, operations and correlations. The transition from quantum to classical is induced by decoherence. The same figure can also represent sets of quantum algorithms or quantum technologies. The distance $\delta$ between the extremal points of both sets along the vertical axis $z$, denoting a selected utility function, represents the quantum advantage.
@@
The analysis presented for the case of complete decoherence highlights the role of Choi–Jamiołkowski isomorphism to analyze geometric features of quantum correlations and channel nonlocality. The standard form of the isomorphism relates bipartite states with quantum operations, while its generalized version links four-partite quantum states with bipartite quantum channels. Due to decoherence the latter suffer a transition to bipartite stochastic matrices ${\cal S}_{d^{2}}$, which describe correlations between both systems. Therefore, the overall geometric structure of the sets of classical and quantum states, channels and correlations is rooted in the geometric structure of the set ${\cal M}_{d^{4}}$ four-partite states (Fig. 5).
Any generalized quantum theory can be recovered from basic principles that force the corresponding set of states to be convex [88]. Hence, if one looks into the set of density matrices from a large perspective, one realizes that it forms a convex set [89]. A closer look reveals its convex subset of separable states, defined as classical mixtures of product states – see Fig. 7. A related structure is characteristic also to the space of quantum maps and quantum correlations. A similar picture can also be used to describe the sets of quantum algorithms and quantum technologies and to visualize the notion of a quantum advantage: The difference between the utility function describing, for instance, efficiency of a given protocol, optimized over all quantum and classical scenarios.
Acknowledgements:   We are thankful to Remigiusz Augusiak, Francesco Buscemi, Dariusz Chruściński, Felix Huber, Kamil Korzekwa, Miguel Navascués, Roberto Salazar, Anna Sanpera and Mario Ziman for discussions and for pointing to references. Support by the Foundation for Polish Science through TEAM-NET project POIR.04.04.00-00-17C1/18-00 and by NCN QuantERA Project No. 2021/03/Y/ST2/00193 is gratefully acknowledged. FS acknowledges projects DESCOM VEGA 2/0183/2 and DeQHOST APVV-22-0570.
@@ Appendix A Proof of Proposition 3
$J^{\prime}_{{A^{\prime}}_{0}^{c}{B^{\prime}}_{0}^{c}A^{\prime}_{1}B^{\prime}_{1}}=\sum_{a,b}$ $(U^{a}_{{R^{\prime}}_{0}^{c}}\otimes V^{b}_{{S^{\prime}}_{0}^{c}}\otimes\mathbf{1}_{{A^{\prime}}_{0}^{c}{B^{\prime}}_{0}^{c}}).$ (53)
$\mathrm{Tr}_{\stackrel{{\scriptstyle\scriptstyle A_{1}R_{0}}}{{B_{1}S_{0}}}}\Big{[}\mathbf{1}_{A_{0}^{c}R_{0}^{c}}\otimes\overline{Z}^{ab}_{\stackrel{{\scriptstyle\scriptstyle A_{1}R_{0}}}{{B_{1}S_{0}}}}\otimes\mathbf{1}_{B_{0}^{c}S_{0}^{c}}\cdot J_{A_{1}B_{1}A_{0}^{c}B_{0}^{c}}\otimes\phi^{+}_{R_{0}R_{0}^{c}}\otimes\phi^{+}_{S_{0}S_{0}^{c}}\Big{]}.(U^{a}_{{R^{\prime}}_{0}^{c}}\otimes V^{b}_{{S^{\prime}}_{0}^{c}}\otimes\mathbf{1}_{{A^{\prime}}_{0}^{c}{B^{\prime}}_{0}^{c}})^{\dagger},$
@@ Appendix B Proof of Proposition 7
Let us show first item (ii). Let $p(a,b|x,y)=\mathrm{Tr}(\sigma P^{a|x}\otimes Q^{b|y})$ determine a quantum probability distribution, where $\sigma$ is a quantum state on $RS$, $\{P^{a|x}\}_{a}$ projective measurements over $R$ and $\{Q^{b|y}\}_{b}$ over $S$. Going to a Stinespring’s dilation, there exist unitary maps $U_{x}:AR\rightarrow AR$ and $V_{y}:BS\rightarrow BS$ such that

### v3

@@ 5 Conclusions and open questions
We infer nonlocal properties of bipartite quantum channels based on the bipartite stochastic matrices produced by them under measurement protocols. In particular, our approach allows us to bound the sets of local operations assisted by shared randomness (LOSR), local operations assisted by shared entanglement (LOSE) and quantum nonsignaling channels (QNS), from the conditional probabilities of local measurement outcomes. This framework is suitable to experimental verification of nonlocality of bipartite channels from local measurements, even in the presence of complete decoherence. In contrast, we show that both entanglement and nonlocality of bipartite quantum channels can be enhanced under nonlocal gradual decoherence.
@@
Figure 7: First glimpse at the structure of multidimensional convex sets of quantum $(\mathcal{Q})$ and classical $(\mathcal{C})$: states, operations and correlations. The transition from quantum to classical is induced by decoherence. The same figure can also represent sets of quantum algorithms or quantum technologies. The distance $\delta$ between the extremal points of both sets along the vertical axis $z$, denoting a selected utility function, represents the quantum advantage.
@@
The analysis presented for the case of complete decoherence highlights the role of Choi–Jamiołkowski isomorphism to analyse geometric features of quantum correlations and channel nonlocality. The standard form of the isomorphism relates bipartite states with quantum operations, while its generalized version links four-partite quantum states with bipartite quantum channels. Due to decoherence the latter suffer a transition to bipartite stochastic matrices ${\cal S}_{d^{2}}$, which describe correlations between both systems. Therefore, the overall geometric structure of the sets of classical and quantum states, channels and correlations is rooted in the geometric structure of the set ${\cal M}_{d^{4}}$ four-partite states (Fig. 5).
Any generalized quantum theory can be recovered from basic principles that force the corresponding set of states to be convex [91]. Hence, if one looks into the set of density matrices from a large perspective, one realizes that it forms a convex set [92]. A closer look reveals its convex subset of separable states, defined as classical mixtures of product states – see Fig. 7. A related structure is characteristic also to the space of quantum maps and quantum correlations. A similar schematic picture can also be used to describe the sets of quantum algorithms and quantum technologies and to visualize the notion of a quantum advantage: the difference between the utility function describing, for instance, efficiency of a given protocol, optimized over all quantum and classical scenarios.
Acknowledgements:   We are thankful to Remigiusz Augusiak, Francesco Buscemi, Dariusz Chruściński, Jakub Czartowski, Felix Huber, Kamil Korzekwa, Miguel Navascués, Roberto Salazar, Anna Sanpera and Mario Ziman for discussions and for pointing to references. Support by the Foundation for Polish Science through TEAM-NET project POIR.04.04.00-00-17C1/18-00 and by NCN QuantERA Project No. 2021/03/Y/ST2/00193 and ERC Advanced Grant TAtypic, project number 101142236 is gratefully acknowledged. FS acknowledges projects DESCOM VEGA 2/0183/2 and DeQHOST APVV-22-0570.
@@ Appendix A Proof of Proposition 3
$J^{\prime}_{{A^{\prime}}_{0}^{c}{B^{\prime}}_{0}^{c}A^{\prime}_{1}B^{\prime}_{1}}=\sum_{a,b}$ $(U^{a}_{{R^{\prime}}_{0}^{c}}\otimes V^{b}_{{S^{\prime}}_{0}^{c}}\otimes\mathbf{1}_{{A^{\prime}}_{0}^{c}{B^{\prime}}_{0}^{c}})$ (53)
$\mathrm{Tr}_{\stackrel{{\scriptstyle\scriptstyle A_{1}R_{0}}}{{B_{1}S_{0}}}}\Big{[}\mathbf{1}_{A_{0}^{c}R_{0}^{c}}\otimes\overline{Z}^{ab}_{\stackrel{{\scriptstyle\scriptstyle A_{1}R_{0}}}{{B_{1}S_{0}}}}\otimes\mathbf{1}_{B_{0}^{c}S_{0}^{c}}\cdot J_{A_{1}B_{1}A_{0}^{c}B_{0}^{c}}\otimes\phi^{+}_{R_{0}R_{0}^{c}}\otimes\phi^{+}_{S_{0}S_{0}^{c}}\Big{]}(U^{a}_{{R^{\prime}}_{0}^{c}}\otimes V^{b}_{{S^{\prime}}_{0}^{c}}\otimes\mathbf{1}_{{A^{\prime}}_{0}^{c}{B^{\prime}}_{0}^{c}})^{\dagger},$
@@ Appendix B Proof of Proposition 7
First we show the second item (ii). Let $p(a,b|x,y)=\mathrm{Tr}(\sigma P^{a|x}\otimes Q^{b|y})$ determine a quantum probability distribution, where $\sigma$ is a quantum state on $RS$, $\{P^{a|x}\}_{a}$ projective measurements over $R$ and $\{Q^{b|y}\}_{b}$ over $S$. Going to a Stinespring’s dilation [93, Theorem 6.9], there exist unitary maps $U_{x}:AR\rightarrow AR$ and $V_{y}:BS\rightarrow BS$ such that